OpenStudy (anonymous):

WILL GIVE MEDAL!!! Simplify the radical expression. sqrt7 +3 sqrt7

OpenStudy (anonymous):

\[A. 2 \sqrt{7} \] \[B. 4\sqrt{14}\] \[C.4\sqrt{7}\] \[D. 2\sqrt{14}\]